Economics
1. Consider the following production function:
Y = F(K, L) = A(2K + 3L)
Does this production function exhibit constant returns to scale?
Answer #1

Y = F (K,L) = A(2K + 3L)

Y = F(XK, XL) = A(2XK + 3XL)

Y = F(XK, XL) = AX(2K + 3L)

As Y = F(XK, XL) = X * F (K,L), this exhibits constant returns to scale.

If this were Y = F(XK, XL) = Xn * F (K,L), where n is > 1, it would have increasing returns to scale or n < 1, it would have decreasing returns returns to scale. Here n = 1 implies constant returns to scale.